In a population of flowers, red is dominant to white. Assuming the population is in HWE, if the red allele frequency is 0.46, what is the red phenotype frequency? answer to two decimals
Answer #1

Frequency of red allele (p) = 0.46

According to HWE

p+q = 1

q = 1-p = 1-0.46 = 0.54

p2 + 2pq + q2 = 1

p2 + 2pq = frequency of red phenotype

q2 = frequency of white phenotype

Frequency of red phenotype = p2 + 2pq

= (0.46)2 + 2*0.46*0.54

= 0.21 + 0.49 = 0.70